# Measuring gravitational lensing time delays with quantum information processing

### v1

@@ VIII.3 Atmospheric fluctuations
We also discuss the noise in $\Delta t$ and the relative phase generated by atmospheric fluctuations and claim that ground-based telescopes are sufficient for our time-delay measurement. Note that the temporal variation in the refractive index of the atmosphere can be quantified by the atmospheric coherence time Roddier (1981); Kellerer and Tokovinin (2007), which is the timescale over which the wave path varies more than a significant fraction of the wavelength. This quantity is typically related to the wavelength and the wind speed. Note that, in our microlensing case, the two branches of the photon take almost the same path in the atmosphere because they are indistinguishable from Earth. Therefore, the refractive index of the atmosphere is the same for both branches, creating no noise in the measured $\Delta t$, provided that the change of refractive index during $\Delta t$ is sufficiently small. A recent result Osborn et al. (2018) gives an estimate of the atmospheric coherence time ($4.18\,\mathrm{ms}$) for the Very Large Telescope, much longer than the upper limit of our observable time delay ($1\,\mathrm{ms}$). Therefore, we claim that atmospheric fluctuations are irrelevant for the stability of our measurement, and it suffices to use ground-based telescopes, which is a much more economical choice—compared with space telescopes—for achieving larger telescope sizes.
@@ IX Time-delay calibration in telescope arrays
A distributed version of our protocol is naturally suited for synchronizing time delays in telescope arrays (Fig. 8). In that scenario Gottesman et al. (2012); Khabiboulline et al. (2019a, b), there are $N$ sites that observe incoming light. We can consider a narrow-band point source, as is the case with an artificial guide star. Then the problems associated with finite-source size and undersampling are alleviated, and the receiver can operate over a small range of frequencies. Learning the delays between sites with small sample complexity enables faster calibration of the array before interferometry, or the ability to use fainter sources.
@@
In contrast to the setup for gravitational lensing discussed earlier, the incoming light arrives at spatially separated sites. If we had lossless channels, we could bring the paths together and make a measurement in frequency space. However, in practice, channels are lossy with exponential degradation in distance, so we use teleportation Gottesman et al. (2012); Khabiboulline et al. (2019a) to overcome this limit, given that entanglement can be purified. Teleportation with minimal entanglement resources necessitates throwing out unnecessary information; hence, we introduce the qubit discretization. We comment on the possibility of a continuous-variable approach employing two-mode squeezed states containing the same number of ebits, but we do not explore this alternative here.
@@ X Summary and discussions
We also emphasize that time-delay measurements in the photon-starved regime may find other applications beyond microlensing delay estimation. In fact, we presented one concrete example for long-baseline quantum telescope arrays Khabiboulline et al. (2019b); Gottesman et al. (2012). These high-resolution interferometers benefit from the capability to determine $\Delta t$ with $t_{c}$ precision from a very limited number of photons, since $\Delta t$ is rapidly changing, the search space is large, or the source is faint. Our approach provides a sample-optimal solution to this problem. Also, since a pointlike narrow-band artificial guide star can be used as the source in this case, the finite-source effect can be avoided, and the physical realization of the frequency-resolving device is much easier.
@@ Appendix A Beam splitter model for dust extinction
The modeling of dust extinction depends on various properties of the interstellar medium. Exploring the exact composition of the interstellar medium is beyond the scope of this paper and is unlikely to succeed due to the limited amount of decisive research progress in this area. However, since we would be satisfied with solid evidence supporting either model, it suffices to study some reasonable approximations of Mie scattering, the general theory of interaction between light and dust particles, and derive upper/lower bounds for the fraction of “bad” photons. Therefore, in this analysis, we consider all dust particles as opaque spherical objects with extinction cross-section $A_{c}$, where the exact value of $A_{c}$ is determined by the Mie theory using particles’ (linear) size $a$ and the wavelength of light $\lambda$ Mie (1908); Li (2008), which is typically upper bounded by a constant factor times the geometric cross section $\pi a^{2}$. When a photon falls within the extinction cross-section of the particle, it is either scattered in various directions or absorbed by the particle. Since the probability that the scattered photon still goes to our telescope is tiny, we assume the photon is lost whenever it is scattered. Under this assumption, we can derive a theory to unify both the random wall model and the beam splitter model as follows.
@@
$g(\eta,\sigma)=\int_{-1}^{1}\eta\sqrt{1-\eta^{2}}p_{\mathrm{var}}(\eta)d\eta=\frac{1}{2}-\sigma^{2}+O(\sigma^{4}).$ (121)
@@
Therefore, to decide which model is closer to reality, we compute the variance of unblocked trapezoids. As mentioned in the above observations, every trapezoid in the final layer is connected to the photon source through a chain of trapezoids with one trapezoid in each layer. This suggests that there is a tree structure in the trapezoids: we let the triangle in the first layer and all trapezoids be nodes in the graph, and let an edge connect two nodes if they are in two adjacent layers and share a base. One can see that there is a unique path from the root node to any other node, hence it is a tree. Moreover, it is a binary tree with $1+\log(d/r)$ layers, and the leaf nodes (nodes in the deepest layer) correspond to the trapezoids in the final layer of the trapezoid model. Observe that the dust particle sampling process is the same as a coloring process where each node in layer $k$ is colored as green with probability $q_{k}$ or red with probability $1-q_{k}$, individually and independently. The event that one trapezoid (with index $i$, for instance) in the final layer is not blocked corresponds to the event that the coloring of the binary tree has an all-green path from the root node to the $i$th leaf node. Now, the number of such all-green paths from the top to the bottom is the number of unblocked trapezoids in the last layer, and the ratio of the average number of such events to the total number of events is also determined by Eqs. (117, 118).
@@
Observe that the variance in the fraction of unblocked trapezoids is approximately inversely proportional to $d/r$, the ratio between the telescope size and the dust particle cross-section. Therefore, the variance is large only if the dust “particles” are so large that their size $r$ is comparable to the telescope aperture $d$, and the “particles” are now huge rocks blocking almost all rays in the triangle. Fortunately, $r\lesssim 1\,\mathrm{\mu m}$ Mathis et al. (1977), hence, for $d\approx 10\,\mathrm{m}$, this ratio is $10^{7}$. Moreover, in our 3D universe, the number of 3D trapezoids (frustums) in the final layer is $d^{2}/r^{2}$ rather than $d/r$, suggesting that the above variance should be replaced by
@@
This implies that, for a typical rate of obtaining a photon in the Baade window $1-p_{\mathrm{loss}}\geq 0.1$ given by Ref. Saha et al. (2019) and Section VII, the standard deviation is $\sim 10^{-7}$, and the probability of any relative variation greater than $5\%$ is upper bounded by $10^{-9}$. Therefore, the fraction of every photon’s wave function that arrives at the telescope is rather stable, and we conclude that the beam splitter model is much closer to the physical reality. This demonstrates the robustness of $\Delta t$ signal against dust extinction.

### v2

@@ VIII.3 Atmospheric fluctuations
We also discuss the noise in $\Delta t$ and the relative phase generated by atmospheric fluctuations and claim that ground-based telescopes are sufficient for our time-delay measurement. Note that the temporal variation in the refractive index of the atmosphere can be quantified by the atmospheric coherence time [97, 98], which is the timescale over which the wave path varies more than a significant fraction of the wavelength. This quantity is typically related to the wavelength and the wind speed. Note that, in our microlensing case, the two branches of the photon take almost the same path in the atmosphere because they are indistinguishable from Earth. Therefore, the refractive index of the atmosphere is the same for both branches, creating no noise in the measured $\Delta t$, provided that the change of refractive index during $\Delta t$ is sufficiently small. A recent result [99] gives an estimate of the atmospheric coherence time ($4.18\,\mathrm{ms}$) for the Very Large Telescope, much longer than the upper limit of our observable time delay ($1\,\mathrm{ms}$). Therefore, we claim that atmospheric fluctuations are irrelevant for the stability of our measurement, and it suffices to use ground-based telescopes, which is a much more economical choice—compared with space telescopes—for achieving larger telescope sizes.
@@ IX Time-delay calibration in telescope arrays
A distributed version of our protocol is naturally suited for synchronizing time delays in telescope arrays (Fig. 7). In that scenario [4, 5, 6], there are $N$ sites that observe incoming light. We can consider a narrow-band point source, as is the case with an artificial guide star. Then the problems associated with finite-source size and undersampling are alleviated, and the receiver can operate over a small range of frequencies. Learning the delays between sites with small sample complexity enables faster calibration of the array before interferometry, or the ability to use fainter sources.
@@
In contrast to the setup for gravitational lensing discussed earlier, the incoming light arrives at spatially separated sites. If we had lossless channels, we could bring the paths together and make a measurement in frequency space. However, in practice, channels are lossy with exponential degradation in distance, so we use teleportation [4, 5] to overcome this limit, given that entanglement can be purified. Teleportation with minimal entanglement resources necessitates throwing out unnecessary information; hence, we introduce the qubit discretization. We comment on the possibility of a continuous-variable approach employing two-mode squeezed states containing the same number of ebits, but we do not explore this alternative here.
@@ X Summary and discussions
We also emphasize that time-delay measurements in the photon-starved regime may find other applications beyond microlensing delay estimation. In fact, we presented one concrete example for long-baseline quantum telescope arrays [6, 4]. These high-resolution interferometers benefit from the capability to determine $\Delta t$ with $t_{c}$ precision from a very limited number of photons, since $\Delta t$ is rapidly changing, the search space is large, or the source is faint. Our approach provides a sample-optimal solution to this problem. Also, since a pointlike narrow-band artificial guide star can be used as the source in this case, the finite-source effect can be avoided, and the physical realization of the frequency-resolving device is much easier.
@@ Appendix A Beam splitter model for dust extinction
The modeling of dust extinction depends on various properties of the interstellar medium. Exploring the exact composition of the interstellar medium is beyond the scope of this paper and is unlikely to succeed due to the limited amount of decisive research progress in this area. However, since we would be satisfied with solid evidence supporting either model, it suffices to study some reasonable approximations of Mie scattering, the general theory of interaction between light and dust particles, and derive upper/lower bounds for the fraction of “bad” photons. Therefore, in this analysis, we consider all dust particles as opaque spherical objects with extinction cross-section $A_{c}$, where the exact value of $A_{c}$ is determined by the Mie theory using particles’ (linear) size $a$ and the wavelength of light $\lambda$ [102, 103], which is typically upper bounded by a constant factor times the geometric cross section $\pi a^{2}$. When a photon falls within the extinction cross-section of the particle, it is either scattered in various directions or absorbed by the particle. Since the probability that the scattered photon still goes to our telescope is tiny, we assume the photon is lost whenever it is scattered. Under this assumption, we can derive a theory to unify both the random wall model and the beam splitter model as follows.
@@
$g(\eta,\sigma)=\int_{-1}^{1}\eta\sqrt{1-\eta^{2}}\cdot p_{\mathrm{var}}(\eta)d\eta=\frac{1}{2}-\sigma^{2}+O(\sigma^{4}).$ (115)
@@
Therefore, to decide which model is closer to reality, we compute the variance of unblocked trapezoids. As mentioned in the above observations, every trapezoid in the final layer is connected to the photon source through a chain of trapezoids with one trapezoid in each layer. This suggests that there is a tree structure in the trapezoids: we let the triangle in the first layer and all trapezoids be nodes in the graph, and let an edge connect two nodes if they are in two adjacent layers and share a base. One can see that there is a unique path from the root node to any other node, hence it is a tree. Moreover, it is a binary tree with $1+\log(d/r)$ layers, and the leaf nodes (nodes in the deepest layer) correspond to the trapezoids in the final layer of the trapezoid model. Observe that the dust particle sampling process is the same as a coloring process where each node in layer $k$ is colored as green with probability $q_{k}$ or red with probability $1-q_{k}$, individually and independently. The event that one trapezoid (with index $i$, for instance) in the final layer is not blocked corresponds to the event that the coloring of the binary tree has an all-green path from the root node to the $i$th leaf node. Now, the number of such all-green paths from the top to the bottom is the number of unblocked trapezoids in the last layer, and the ratio of the average number of such events to the total number of events is also determined by Eqs. (111, 112).
@@
Observe that the variance in the fraction of unblocked trapezoids is approximately inversely proportional to $d/r$, the ratio between the telescope size and the dust particle cross-section. Therefore, the variance is large only if the dust “particles” are so large that their size $r$ is comparable to the telescope aperture $d$, and the “particles” are now huge rocks blocking almost all rays in the triangle. Fortunately, $r\lesssim 1\,\mathrm{\mu m}$ [93], hence, for $d\approx 10\,\mathrm{m}$, this ratio is $10^{7}$. Moreover, in our 3D universe, the number of 3D trapezoids (frustums) in the final layer is $d^{2}/r^{2}$ rather than $d/r$, suggesting that the above variance should be replaced by
@@
This implies that, for a typical rate of obtaining a photon in the Baade window $1-p_{\mathrm{loss}}\geq 0.1$ given by Ref. [87] and Section VII, the standard deviation is $\sim 10^{-7}$, and the probability of any relative variation greater than $5\,\%$ is upper bounded by $10^{-9}$. Therefore, the fraction of every photon’s wave function that arrives at the telescope is rather stable, and we conclude that the beam splitter model is much closer to the physical reality. This demonstrates the robustness of $\Delta t$ signal against dust extinction.
Appendix B Narrowband flare combination
In this appendix, we discuss the flare combination approach when the coherence time $t_{c}$ is much longer than the difference in $\Delta t$ between different flares, i.e., when the photons are from a narrow-band source. In this case, peaks in the score function corresponding to different $\Delta t_{i}$s overlap with each other, so the data processing is much easier than in the broadband case. This may not occur in practice because the emissions of stellar flares are unlikely to be restricted within several $\mathrm{GHz}$. Nonetheless, this regime is potentially relevant to Section IX, when multiple guidestars are used, and artificial guidestars could be narrowband photon sources.
We introduce a new score function for the $i$th flare, denoted by $f_{\mathrm{c},i}$, and a score function for the combination of all $m$ flares, denoted by $G_{\mathrm{\Delta t_{0}}}$:
$\mathbb{E}[\exp(\mathrm{i}\nu_{i,j}\tau)]$ $=\int_{-\infty}^{\infty}e^{\mathrm{i}\nu_{i,j}\tau}p_{A}(\nu_{i,j}|\Delta t)d\nu_{i,j}$ (126)
$=\sum_{j=1}^{n}\mathopen{}\mathclose{{\left(\mathbb{E}[R_{j}^{2}]+\mathbb{E}[I_{j}^{2}]}}\right)=\sum_{j=1}^{n}\mathbb{E}[R_{j}^{2}+I_{j}^{2}]d=n.$
$\mathbb{E}[f_{\mathrm{c},i}]$ $=\mathopen{}\mathclose{{\left|\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{j=1}^{n}\exp(\mathrm{i}\nu_{i,j}\tau)}}\right]}}\right|^{2}+\operatorname{Var}\mathopen{}\mathclose{{\left(\sum_{j=1}^{n}\exp(\mathrm{i}\nu_{i,j}\tau)}}\right)$ (128)
$=\frac{n^{2}Q^{2}\gamma_{A}^{2}}{4}+\sum_{j=1}^{n}\mathbb{E}[R_{j}^{2}+I_{j}^{2}]-\sum_{j=1}^{n}\mathopen{}\mathclose{{\left[\mathbb{E}[R_{j}]^{2}-\mathbb{E}[I_{j}]^{2}}}\right]$
$=\frac{n^{2}Q^{2}\gamma_{A}^{2}}{4}-\frac{nQ\gamma_{A}^{2}}{4}+n$
$=\frac{\gamma_{A}^{2}}{4}(n_{\mathrm{sig}}^{2}-n_{\mathrm{sig}})+n$
[…]
Figure 9: Numerical and analytical results for the flare number $m$ necessary to yield a detection as a function the number of signal photons $n_{\mathrm{sig}}$ per flare under various settings. All curves are the analytically-derived sufficient conditions (see Eq. 135), and points represent the sufficient conditions obtained from numerical simulation, which are closer to optimality. The black curves and points are the baseline setting discussed in our M-dwarf flare observation proposal: $T/t_{c}=10^{4},Q=0.4$, and confidence level is $95\%$. The red curves and points show variations on these three parameters, with one parameter varied in each plot. All simulations are based on magnification $A=1.34$. In (a), we compare the flare number requirement for different confidence levels ($95\%$ and $50\%$). In (b), we alter the signal-to-noise ratio $Q$ from $0.4$ to $0.6$. In (c) we change $T/t_{c}$ from $10^{4}$ to $10^{2}$.
$\quad\Pr[f_{\mathrm{c},i}(\tau)-n>\epsilon]$ (131)
$\leq\exp(-\frac{(\sqrt{\epsilon+n}-\sqrt{n})^{2}}{2n})$
$=\exp(-\frac{\epsilon}{2n}-1+\sqrt{1+\frac{\epsilon}{n}}).$
The above upper bound converges to $\exp(-\frac{\epsilon^{2}}{8n^{2}})$ when $\epsilon\ll n$ and becomes $\exp(-\frac{\epsilon}{2n})$ when $\epsilon\gg n$. This implies that $f_{\mathrm{c},i}$ is approximately a sub-exponential random variable with parameters $(4n^{2},n)$. Therefore, as a sum of sub-exponential random variables, Bernstein’s inequality [104] gives the following bounds for $G_{\mathrm{\Delta t_{0}}}$:
$\quad\Pr[G_{\mathrm{\Delta t_{0}}}(\tau)>G_{\mathrm{th}}]$ (132)
[…]
and for $n_{\mathrm{sig}}$ close to $2$, the sufficient condition is
Again, we emphasize that the above derivation only gives the sufficient condition for a successful time delay measurement, or an upper bound (in fact, a loose upper bound) of $m$ for any given $(n_{\mathrm{sig}},T/t_{c},Q,A)$ parameters. The actual lower bound for $m$ will likely have different constant factors, but will never be higher than the upper bounds given by the sufficient-condition analysis. Indeed, we perform numerical simulations to find the number of flares actually needed for various realistic parameters, and the results are significantly less than the theoretical upper bounds. We present the numerical results and their comparison with theoretical sufficient conditions in Fig. 9. We also give one realistic example for the fiducial M-dwarf flare setting where $A=1.34$ (hence $\gamma_{A}=0.666$) and $Q=0.4$. If we set $T/t_{c}=10000$, use the Extremely Large Telescope (which gives $n_{\mathrm{sig}}=426$), and aim for $95\,\%$ confidence, then the sufficient condition given by our the analytic bound is $m\geq 3$, while numerical simulation shows that even $m=1$ works.
To give an intuitive demonstration of how our algorithm works, in Fig. 10, we present the score function values from the numerical simulation of one successful time-delay measurement. The results are from (the simulation of) the same fiducial observation of M dwarf flares but using Keck rather than the Extremely Large Telescope. It combines 5 flares with 66 signal photons and 103 noise photons per flare.